# Lessons Learned from Developing a Privacy-Preserving Multimodal Wearable for Local Voice-and-Vision Inference

### v1

@@ Abstract.
Many promising applications of multimodal wearables require continuous sensing and heavy computation, yet users reject such devices due to privacy concerns. This paper shares our experiences building an ear-mounted voice-and-vision wearable that performs local AI inference using a paired smartphone as a trusted personal edge. We describe the hardware–software co-design of this privacy-preserving system, including challenges in integrating a camera, microphone, and speaker within a 30-gram form factor, enabling wake word-triggered capture, and running quantized vision-language and large-language models entirely offline. Through iterative prototyping, we identify key design hurdles in power budgeting, connectivity, latency, and social acceptability. Our initial evaluation shows that fully local multimodal inference is feasible on commodity mobile hardware with interactive latency. We conclude with design lessons for researchers developing embedded AI systems that balance privacy, responsiveness, and usability in everyday settings.
@@ 1. Introduction
Wearable devices promise contextual intelligence, but achieving this vision requires always-available sensing and computation—precisely the properties that raise user distrust. Continuous data capture and cloud inference have fueled privacy backlash, leading to the rejection of devices such as camera glasses and body-mounted loggers (Hoyle2014PrivacyBehaviorsLifeloggers). Prior work on privacy-aware lifelogging and visual privacy protection (Korayem2014ScreenAvoider; shu2016cardeacontextawarevisualprivacy) highlights how user discomfort often stems from ambiguity about when recording occurs and how data is used. These lessons remain highly relevant as new wearable devices emerge, embedding microphones, cameras, and sensors near the head for continuous multimodal interaction (Hu2025EarableSurvey; Fan2023DesignEarableSensing; Elfouly2025WearableComputingComprehensiveSurvey).
At the same time, advances in mobile accelerators and compact multimodal models make it increasingly feasible to perform meaningful inference locally, without uploading raw data to external servers. Our work explores this convergence: a lightweight, privacy-preserving form factor that supports large-model reasoning for assistive tasks such as situational awareness (“what am I looking at?”), conversational assistance, and potentially memory support. We examine the system-level challenges of enabling large-model reasoning under the strict constraints of a small, user-facing device—balancing power, latency, comfort, and trust while keeping computation local to the user’s own devices. By demonstrating that local multimodal inference is practical within these limits, this work lays the groundwork for future applications in memory augmentation, contextual recall, and personalized assistive intelligence.
@@ 2. Background and Related Work
Wearable sensing systems have long aimed to provide continuous, context-sensitive intelligence, but early lifelogging work revealed persistent challenges around privacy, social acceptability, and ambiguity in recording practices. Continuous first-person capture raises concerns for both wearers and bystanders, with perceived sensitivity strongly shaped by location, activity, and who or what appears in view (Hoyle2014PrivacyBehaviorsLifeloggers; Yoon2017EthicalUseLifelogging; Sellen2010BeyondTotalRecall). Empirical reviews of deployments in health and daily-life settings catalog practical precautions and residual risks associated with wearable cameras (Meyer2022WearableCamerasEthics; Meyer2023WearableCamerasDailyLife), while recent analyses show how egocentric video can leak personal traits well beyond user intent (Li2025EgoPrivacy; article). Complementary technical defenses include redacting sensitive on-screen content (Korayem2014ScreenAvoider), providing context-aware protections for bystanders (Shu2016Cardea), and mechanisms that allow people or environments to signal opt-out preferences (Raval2016MarkIt; Denning2014SmartphoneCamera). Our design responds to this body of work by defaulting to short, event-driven capture and confining all processing to user-owned devices.
@@
Advances in on-device multimodal inference further reshape the design space. Modern smartphones now include NPUs and GPUs capable of running compact language and vision-language models locally, including Qwen2-VL–style architectures optimized for edge deployment (Qwen2VL2024) and FastVLM-style pipelines that target mobile and embedded platforms (FastVLM2025). Lightweight visual encoders such as FastViT reduce memory and compute requirements for mobile vision workloads (Vasu2023FastViT). Prior work on mobile and embedded ML has also explored software and model-level optimizations, from DeepX and DeepSense for resource-efficient inference on phones and sensor data (Lane2015DeepX; Huynh2017DeepSense) to DeepIoT and integer-only quantization schemes tailored to low-power devices (Yao2017DeepIoT; Jacob2018Quantization). We build on these developments via MLX (mlx2023) and Apple’s on-device ASR framework (AppleSFSpeechRecognizer) to support fully offline speech, language, and vision-language inference, ensuring that all sensor data remains on the user’s smartphone.
For wake word detection, keyword spotting (KWS) on microcontrollers is backed by TinyML literature. The Speech Commands dataset provides a widely used benchmark for limited-vocabulary speech recognition (Warden2018SpeechCommands), while convolutional architectures such as small-footprint CNNs, TC-ResNet variants, and microcontroller-focused models trade accuracy against memory and latency (Sainath2015KWS; Choi2019TCResNet; Zhang2017HelloEdgeKWS). Our wake word model follows this line of work but emphasizes pragmatic choices to close the train–deploy gap, including synthetic-voice augmentation and hardware re-recording on the actual device microphone.
Finally, efficient on-device routing and intent classification is well supported by classic text-categorization techniques. TF–IDF features combined with linear models remain strong baselines for short utterances (Sebastiani2002TextCategorization; Pedregosa2011ScikitLearn). In our system, such lightweight classifiers route queries to device-control, visual, or conversational pipelines, reserving heavier LLM/VLM inference for the final stage.
@@ 3.1. Hardware Integration
The device integrates a small camera (OV5460 AF), microphone (MEMS from MCU) , speaker (1w 8ohm) paired with amp (Class D 3w 8-4ohm), and a Wi-Fi/BLE-capable microcontroller (XIAO ESP32S3 Sense) inside a pen-shaped 3D-printed enclosure that hooks over the ear (Figure 1). The camera near the temple approximates the user’s field of view. A 200 mAh Li-ion battery along the behind-ear segment powers the device. The total weight is approximately 30 g. BLE is used solely for device discovery and secure provisioning (e.g., setting Wi-Fi credentials). After setup, all runtime communication (audio, discrete photo snapshots, and control messages) runs over Wi-Fi, which offers sufficient bandwidth and more predictable latency for interactive inference. We fabricated three functional prototypes with this architecture, experimenting with different components and enclosure materials: PLA, resin, TPU (Figure 1).
Bring-up via a minimal webserver. A turning point in the hardware-software co-design was treating the MCU as a ”pen-webserver” during the bring-up. Before we attempted any wake word detection or multimodal streaming, we deployed a minimal TCP web socket server firmware which was derived from the CameraWebServer example in the XIAO ESP32S3 sense documentation. We exposed endpoints for toggling an on-board LED, capturing a single camera frame, and sending/receiving short audio buffers.
We accelerate early hardware–software integration by first treating the MCU as a simple ‘pen web server”.’ This minimal firmware (with endpoints to blink an LED, snap a photo, stream audio, etc.) let us quickly debug power, connectivity, and sensor functionality via simple web requests (e.g., using curl from a laptop). This approach forced us to settle critical details early (IP advertising, Wi-Fi reconnection after sleep, control message formats) before adding the full inference pipeline.
@@
Bringing up this pipeline required several iterations. Early unidirectional streams revealed clicks and pops caused by encoding artifacts and buffer under-runs; switching to a ring buffer with $\sim$300 ms of play-out delay eliminated these issues while keeping latency acceptable. Adding the reverse audio path exposed new failures—repeated samples, degraded playback, and occasional stalls—that initially appeared to be decoder bugs. Instead, they stemmed from Wi-Fi/BLE coexistence: both radios share the 2.4 GHz path, and bidirectional audio stressed the automatic coexistence policy. Explicitly prioritizing BLE only during provisioning and Wi-Fi during normal operation resolved these instabilities without hardware changes.
@@ 3.2. Power Budgeting and Battery Life
Our ear-mounted prototype uses a 200 mAh Li-ion cell to power a XIAO ESP32S3 Sense, an always-on microphone, Wi-Fi/BLE radios, and a 3 W class-D amplifier driving a 1 W speaker. Board measurements and vendor data indicate that light-sleep draws only a few milliamps, modem-sleep with the radio ready consumes on the order of 25 mA, and deep-sleep falls to tens of microamps. Recording audio increases the board current to roughly 50–60 mA, while enabling Wi-Fi or BLE in active mode pushes consumption to about 80–100 mA. The on-board camera is particularly expensive: capturing VGA-resolution frames can add hundreds of milliamps in short bursts. A small class-D audio amplifier contributes only a few milliamps of quiescent current when idle, but during 1 W playback its input power corresponds to roughly 300 mA at 4 V.
@@
Baseline (listening for wake word) MCU in modem-sleep, mic + MFCC pipeline armed for wake word, BLE advertising at low duty cycle, amplifier in standby $\approx$80–100 mA $\approx$2 h
Active query burst Wi-Fi TX/RX, camera capture, mic streaming, amplifier playing back speech, control logic active $\approx$400–450 mA $\approx$25–30 min (if continuous)
In practice, active queries last only a few seconds, so these high-power episodes contribute relatively little energy compared to the cost of keeping the wake word path armed. If the device were to run in “always listening” mode continuously, the 200 mAh cell would be exhausted after roughly two hours. To approach day-scale usage, the system must therefore spend most of its time in deep or light sleep and only enable wake word listening in short, user-initiated sessions (for example, a 30-minute walking or commuting interval). This trade-off small, comfortable battery versus continuous availability ultimately led us to design the current prototype for short interaction sessions rather than true 24-hour standby.
@@ 3.3. Local Inference Pipeline
Triggered by a wake word, the device records the user’s query audio and, for vision-grounded queries, triggers a single photo capture. Both are sent over Wi-Fi to the paired smartphone, which executes the entire inference pipeline locally using Core ML and MLX.
Figure 3 shows the overall inference flow.
@@
Conversational queries run on a locally hosted LLaMA 3.2 1B model, quantized to 4-bit and served via MLX with Metal GPU acceleration. Inference is gated on foreground app state to avoid iOS background GPU issues; GPU caches are cleared between calls, using 128–384 MB depending on device.
@@
The pipeline operates without cloud calls or third-party servers. Adaptive GPU cache sizing, lazy model loading, and compressed audio streaming keep latency low while respecting memory constraints. Internal tests on an iPhone 14 show end-to-end latency of roughly 2–3 seconds for typical queries.
@@ 3.4. Intelligence Architecture
Incoming user queries—typically transcribed speech from the wake word pipeline— first pass through a lightweight intent classifier implemented using a TF–IDF + Logistic Regression model. This model runs entirely on-device and routes text into one of four categories: device_control, visual_query, general_question, or conversational. The classifier executes within tens of milliseconds and determines which downstream processing path to trigger: direct device commands (e.g., “take a photo”), visual queries invoking the FastVLM model, or conversational responses handled by the LLaMA 3.2 1B language model. This simple text-based classifier serves as the entry point of the system’s intelligence hierarchy and minimizes latency by filtering non-LLM interactions early.
@@
The intent router manages pipeline execution after classification. It orchestrates three main pathways: (1) device control, which issues commands directly to the pen firmware; (2) visual queries, which perform image capture followed by FastVLM analysis; and (3) conversational queries, which invoke the language model with optional memory retrieval. The router enforces strict foreground execution and model loading to uphold privacy guarantees.
@@
The heavier models in the pipeline—LLaMA 3.2 1B (quantized to 4-bit) for text and FastVLM (Qwen2-VL) for vision-language understanding— are loaded dynamically only while the app is in the foreground. Inference is performed fully locally using Apple’s MLX and Core ML frameworks, ensuring no data leaves the user’s device. Intermediate results (transcripts, embeddings, and scene descriptors) are stored in encrypted Core Data containers, and the app maintains a zero-API-call policy for all intelligence functions.
@@ 3.5. Wake word Dataset Design
Reliable wake word detection on a small MCU required a dataset that captured how people might naturally say the phrase “hey dotty” across voices, speaking styles, and environments, while also covering common confusions. To bootstrap this dataset without recruiting large numbers of speakers, we used a commercial neural text-to-speech (TTS) system to synthesize diverse, high-quality training utterances.
Synthetic wake word positives. We curated roughly 50 synthetic voices (roughly balanced across perceived male and female timbres) and manually filtered out any that sounded obviously robotic. For each retained voice we generated multiple renditions of the wake word “hey dotty”. These renditions varied pitch (from higher to lower pitched voice), speaking rate (0.6-1.0 s duration), and prosody (different intonations/deliveries) across synthesized utterances.
@@
Re-recording through the device. Rather than train directly on clean TTS waveforms, we played each generated utterance through a mono speaker and re-recorded it with the actual pen hardware. The pen was positioned and oriented as in the CAD model microphone facing away from the speaker, on the “far side” of a head mock-up so the captured signal reflected the real acoustic path and microphone characteristics. We repeated this in several ambient environments (e.g., quiet room, office-style background chatter, white-noise ventilation, cafe-like ambiance, and other playlists of ambient noise) as well as in near-silence. For each voice and environment we aimed to collect multiple clean renditions of the wake word, yielding on the order of 20 usable samples per voice and a total of several hundred short clips ($\approx 7$ minutes of labeled audio).
Negative and confuser classes. To reduce false positives from phonetically similar phrases, we created a dedicated confuser class containing utterances such as “hey daddy,” “hey dobby,” and other near-rhymes. Additional random negatives consisted of short arbitrary phrases and clipped fragments (e.g., only “hey d-” or “-dotty”) that resulted when the wake word straddled a recording boundary. Finally, we added a noise class of environment-only audio with no foreground speech. Collecting these classes via the same re-recording setup ensured that microphone response and room acoustics were consistent across positives and negatives.
Iterating on segmentation. Our initial collection script stored one-second WAV files to an SD card and required the speaker to say the word inside each interval. In practice this led to many clips where the word was partially cut at the start or end. Midway through development we switched to continuous recording and performed segmentation offline using a waveform viewer (via the training tool): we visually identified energy spikes corresponding to the wake word, previewed them, and exported tightly trimmed segments. This change significantly improved dataset quality at the cost of some manual curation.
Data organization. All final clips were stored as WAV files organized into label-based folders (heydotty, confuse, noise, unknown/random) to match the expected input format of our training pipeline. This structure also simplified later experimentation with class rebalancing (e.g., oversampling confusers to further reduce false triggers). Although we ultimately used only a subset of the available synthetic voices ($\approx 10$ male, 7 female) due to time constraints, the process is easily extensible to a larger pool in future iterations.

### v2

@@ Abstract.
Many promising applications of multimodal wearables require continuous sensing and heavy computation, yet users reject such devices due to privacy concerns. This paper shares our experiences building an ear-mounted voice-and-vision wearable that performs local AI inference using a paired smartphone as a trusted personal edge. We describe the hardware-software co-design of this privacy-preserving system, including challenges in integrating a camera, microphone, and speaker within a 30-gram form factor, enabling wake word-triggered capture, and running quantized vision-language and large-language models entirely offline. Through iterative prototyping, we identify design hurdles in power budgeting, connectivity, latency, and social acceptability. Our initial evaluation shows that fully local multimodal inference is feasible on commodity mobile hardware with interactive latency. We conclude with design lessons for researchers developing embedded AI systems that balance privacy, responsiveness, and usability in everyday settings.
@@ 1. Introduction
Wearable devices promise contextual intelligence, but achieving this vision requires always-available sensing and computation. Continuous data capture and cloud inference have fueled privacy backlash, leading to the rejection of devices such as camera glasses and body-mounted loggers (Hoyle2014PrivacyBehaviorsLifeloggers). Prior work on privacy-aware lifelogging and visual privacy protection (Korayem2014ScreenAvoider; shu2016cardeacontextawarevisualprivacy) highlights how user discomfort often stems from ambiguity about when recording occurs and how data is used. These lessons remain highly relevant as new wearable devices emerge, embedding microphones, cameras, and sensors near the head for continuous multimodal interaction (Hu2025EarableSurvey; Fan2023DesignEarableSensing; Elfouly2025WearableComputingComprehensiveSurvey).
At the same time, advances in mobile accelerators and compact multimodal models make it increasingly feasible to perform meaningful inference locally, without uploading raw data to external servers. Our work explores this convergence: a lightweight, privacy-preserving form factor that supports large-model reasoning for assistive tasks such as situational awareness (“what am I looking at?”), conversational assistance, and potentially memory support. We examine the system-level challenges of enabling large-model reasoning under the strict constraints of a small, user-facing device—balancing power, latency, comfort, and trust while keeping computation local to the user’s own devices. By demonstrating that local multimodal inference is practical within these limits, this work lays the groundwork for applications in memory augmentation, contextual recall, and personalized assistive intelligence.
@@ 2. Background and Related Work
Wearable sensing systems have long aimed to provide continuous, context-sensitive intelligence, but early lifelogging work revealed persistent challenges around privacy, social acceptability, and ambiguity in recording practices. Continuous first-person capture raises concerns for both wearers and bystanders, with perceived sensitivity strongly shaped by location, activity, and who or what appears in view (Hoyle2014PrivacyBehaviorsLifeloggers; Yoon2017EthicalUseLifelogging; Sellen2010BeyondTotalRecall). Empirical reviews of deployments in health and everyday settings catalog practical precautions and residual risks associated with wearable cameras (Meyer2022WearableCamerasEthics; Meyer2023WearableCamerasDailyLife), while recent analyses show how egocentric video can leak personal traits well beyond user intent (Li2025EgoPrivacy; article). Complementary technical defenses include redacting sensitive on-screen content (Korayem2014ScreenAvoider), providing context-aware protections for bystanders (Shu2016Cardea), and mechanisms that allow people or environments to signal opt-out preferences (Raval2016MarkIt; Denning2014SmartphoneCamera). Our design responds to this body of work by defaulting to short, event-driven capture and confining all processing to user-owned devices.
@@
Advances in on-device multimodal inference further reshape the design space. Modern smartphones now include NPUs and GPUs capable of running compact language and vision-language models locally, including Qwen2-VL–style architectures optimized for edge deployment (Qwen2VL2024) and FastVLM-style pipelines that target mobile and embedded platforms (FastVLM2025). Lightweight visual encoders such as FastViT reduce memory and compute requirements for mobile vision workloads (Vasu2023FastViT). Prior work on mobile and embedded ML has also explored software and model-level optimizations, from DeepX and DeepSense for resource-efficient inference on phones and sensor data (Lane2015DeepX; Huynh2017DeepSense) to DeepIoT and integer-only quantization schemes tailored to low-power devices (Yao2017DeepIoT; Jacob2018Quantization). We build on these developments via MLX (mlx2023) and Apple’s on-device ASR framework (AppleSFSpeechRecognizer) to support fully offline speech, language, and vision-language inference, with all sensor data remaining on the user’s smartphone.
For wake word detection, keyword spotting (KWS) on microcontrollers is backed by TinyML literature. The Speech Commands dataset provides a widely used benchmark for limited-vocabulary speech recognition (Warden2018SpeechCommands), while convolutional architectures such as small-footprint CNNs, TC-ResNet variants, and microcontroller-focused models trade accuracy against memory and latency (Sainath2015KWS; Choi2019TCResNet; Zhang2017HelloEdgeKWS). Our wake word model follows this line of work but emphasizes pragmatic choices to close the training–deployment gap, including synthetic-voice augmentation and hardware re-recording on the actual device microphone.
Finally, efficient on-device routing and intent classification is well supported by classic text-categorization techniques. TF–IDF features combined with linear models remain strong baselines for short utterances (Sebastiani2002TextCategorization; Pedregosa2011ScikitLearn). In our system, such lightweight classifiers route queries to device control, visual, or conversational pipelines, reserving heavier LLM/VLM inference for the final stage.
@@ 3.1. Hardware Integration
The device integrates a small camera (OV5460 AF), microphone (MEMS from MCU) , speaker (1W, $8\Omega$) paired with amp (Class D, 3W, $8-4\Omega$), and a Wi-Fi/BLE-capable microcontroller (XIAO ESP32S3 Sense) inside a 3D-printed enclosure that hooks over the ear (Figure 1). The camera near the temple approximates the user’s field of view. A 200 mAh Li-ion battery along the behind-ear segment powers the device. The total weight is approximately 30 g. BLE is used solely for device discovery and secure provisioning (e.g., setting Wi-Fi credentials). After setup, all runtime communication (audio, discrete photo snapshots, and control messages) runs over Wi-Fi/mobile hotspot, which offers sufficient bandwidth and more predictable latency for interactive inference. We fabricated three functional prototypes with this architecture, experimenting with different components and enclosure materials: PLA, resin, TPU (Figure 1).
Bring-up via a minimal webserver. A turning point in the hardware-software co-design was treating the MCU as a standalone webserver during bring-up. Before attempting wake word detection or multimodal streaming, we deployed a minimal TCP WebSocket server derived from the CameraWebServer example in the XIAO ESP32S3 Sense documentation. We exposed endpoints for toggling an on-board LED, capturing a single camera frame, and sending/receiving short audio buffers. Interacting with these endpoints via simple web requests (e.g., curl) let us rapidly validate power behavior, connectivity, and sensor functionality. This also forced us to resolve early system details before layering on the full inference pipeline, including IP advertising, Wi-Fi reconnection after sleep, and control message formats.
@@
Bringing up this pipeline required several iterations. Early unidirectional streams yielded clicks and pops caused by encoding artifacts and buffer under-runs; switching to a ring buffer with $\sim$300 ms of play-out delay eliminated these issues while keeping latency acceptable. Adding the reverse audio path exposed new failures–repeated samples, degraded playback, and occasional stalls–that initially appeared to be decoder bugs. Instead, they stemmed from Wi-Fi/BLE coexistence: both radios share the 2.4 GHz path, and bidirectional audio stressed the automatic coexistence policy. Explicitly prioritizing BLE only during provisioning and Wi-Fi during normal operation resolved these instabilities without hardware changes.
@@ 3.2. Power Budgeting and Battery Life
Our ear-mounted prototype uses a 200 mAh Li-ion cell to power a XIAO ESP32S3 Sense, Wi-Fi/BLE radios, and a 3W Class D amplifier driving a 1W speaker. Board measurements and vendor data indicate that light-sleep draws only a few milliamps, modem-sleep with the radio ready consumes on the order of 25 mA, and deep-sleep falls to tens of microamps. Recording audio increases board current to roughly 50-60 mA, while enabling Wi-Fi or BLE in active mode pushes consumption to about 80-100 mA. The on-board camera is particularly expensive: capturing VGA-resolution frames can add hundreds of milliamps in short bursts. The Class D audio amplifier contributes only a few milliamps of quiescent current when idle, but during 1 W playback its input power corresponds to roughly 300 mA at 4 V.
@@
Baseline (listening for wake word) MCU in modem-sleep, mic + MFCC pipeline armed for wake word, BLE advertising at low duty cycle, amplifier in standby $\approx$80-100 mA $\approx$2 h
Active query burst Wi-Fi TX/RX, camera capture, mic streaming, amplifier playing back speech, control logic active $\approx$400-450 mA $\approx$25-30 min (if continuous)
In practice, active queries last only a few seconds, so these high-power episodes contribute relatively little energy compared to the cost of keeping the wake word path armed. If the device were to run in “always listening” mode continuously, the 200 mAh cell would be exhausted after roughly two hours. To approach day-scale usage, the system must therefore spend most of its time in deep or light sleep and only enable wake word listening in short, user-initiated sessions (for example, a 30-minute walking or commuting interval). This trade-off, using a small, lightweight battery versus continuous availability ultimately led us to design the current prototype for short interaction sessions rather than true 24-hour standby.
@@ 3.3. Local Inference Pipeline
Triggered by a wake word, the device records the user’s query audio and, for vision-grounded queries, triggers a single photo capture. Both are sent over Wi-Fi to the paired smartphone, which executes the entire inference pipeline locally using Core ML and MLX. Figure 3 shows the overall inference flow.
@@
Conversational queries run on a locally hosted LLaMA 3.2 1B model, quantized to 4-bit and served via MLX with Metal GPU acceleration. Inference is gated on foreground app state to avoid iOS background GPU issues; GPU caches are cleared between calls, using 128-384 MB depending on device.
@@
The pipeline operates without cloud calls or third-party servers. Adaptive GPU cache sizing, lazy model loading, and compressed audio streaming keep latency low while respecting memory constraints. Internal tests on an iPhone 14 show end-to-end latency of roughly 2-3 seconds for typical queries.
@@ 3.4. Intelligence Architecture
Incoming user queries—typically transcribed speech from the wake word pipeline— first pass through a lightweight intent classifier implemented using a TF-IDF + Logistic Regression model. This model runs entirely on-device and routes text into one of four categories: device_control, visual_query, general_question, or conversational. The classifier executes within tens of milliseconds and determines which downstream processing path to trigger: direct device commands (e.g., “take a photo”), visual queries invoking the FastVLM model, or conversational responses handled by the LLaMA 3.2 1B language model. This simple text-based classifier serves as the entry point of the system’s intelligence hierarchy and minimizes latency by filtering non-LLM interactions early.
@@
The intent router manages pipeline execution after classification. It orchestrates three main pathways: (1) device control, which issues commands directly to the earpiece firmware; (2) visual queries, which perform image capture followed by FastVLM analysis; and (3) conversational queries, which invoke the language model with optional memory retrieval. The router enforces strict foreground execution and model loading to uphold privacy guarantees.
@@
The heavier models in the pipeline–LLaMA 3.2 1B (quantized to 4-bit) for text and FastVLM (Qwen2-VL) for vision-language understanding–are loaded dynamically only while the app is in the foreground. Inference is performed fully locally using Apple’s MLX and Core ML frameworks, ensuring no data leaves the user’s device. Intermediate results (transcripts, embeddings, and scene descriptors) are stored in encrypted Core Data containers, and the app maintains a zero-API-call policy for all intelligence functions.
@@ 3.5. Wake word Dataset Design
Reliable wake word detection on a small MCU required a dataset that captured how people might naturally say the chosen phrase across voices, speaking styles, and environments, while also covering common confusions. The phrase of choice is arbitrary, and for this prototype we went with ”hey dotty”. To bootstrap this dataset without large numbers of speakers, we used a commercial neural text-to-speech (TTS) system to synthesize diverse, high-quality training utterances.
Synthetic wake word positives. We curated roughly 50 synthetic voices (roughly balanced across perceived male and female timbres) and manually filtered out any that sounded obviously robotic. For each retained voice we generated multiple renditions of the wake word. These renditions varied in pitch, speaking rate (0.6-1.0 s duration), and prosody (different intonations/deliveries) across synthesized utterances.
@@
Re-recording through the device. Rather than train directly on clean TTS waveforms, we played each generated utterance through a mono speaker and re-recorded it with the actual device. The earpiece was worn and positioned with the microphone facing away from the speaker, on the “far side” of a head mock-up so the captured signal reflected the real acoustic path and microphone characteristics. We repeated this in several ambient environments (e.g., quiet room, office-style background chatter, white-noise ventilation, cafe-like ambiance, and other playlists of ambient noise) as well as in near-silence. For each voice and environment we aimed to collect multiple clean renditions of the wake word, yielding on the order of 20 usable samples per voice and a total of several hundred short clips (around 7 minutes of labeled audio).
Negative and confuser classes. To reduce false positives from phonetically similar phrases, we created a dedicated confuser class containing utterances such as “hey dobby,” and other near-rhymes. Additional random negatives consisted of short arbitrary phrases and clipped fragments (e.g., only “hey d-” or “-dotty”) that resulted when the wake word straddled a recording boundary. Finally, we added a noise class of environment-only audio with no foreground speech. Collecting these classes via the same re-recording setup ensured that microphone response and room acoustics were consistent across positives and negatives.
Iterating on segmentation. Our initial collection script stored one-second WAV files to an SD card and required the speaker to say the word inside each interval. In practice this led to many clips where the word was partially cut at the start or end. Midway through development we switched to continuous recording and performed segmentation offline using a waveform viewer: we visually identified energy spikes corresponding to the wake word, previewed them, and exported tightly trimmed segments. This change significantly improved dataset quality at the cost of some manual curation.
Data organization. All final clips were stored as WAV files organized into label-based folders (heydotty, confuse, noise, unknown/random) to match the expected input format of our training pipeline. This structure also simplified later experimentation with class rebalancing (e.g., oversampling confusers to further reduce false triggers). Although we ultimately used only a subset of the available synthetic voices (10 male, 7 female) due to time constraints, the process is easily extensible to a larger pool in future iterations.